OpenStudy (anonymous):

A circle has a radius of 9 ft. What is the approximate area of the circle? Use 3.14 to approximate pi. Write the answer to the nearest hundredth. I am thinking the answer is 254ft squared

OpenStudy (anonymous):

It is 254.47

OpenStudy (anonymous):

if we round to the nearest hundredth would that be 254?

OpenStudy (anonymous):

Formula: \(A = \pi r^2\) Plug in what we know: \(A = (3.14)(9^2)\) Multiply 3.14 * 9 * 9

OpenStudy (anonymous):

@pranavsrivatsan

OpenStudy (anonymous):

Yes it will be, rounded. :)

OpenStudy (anonymous):

@thedomainbunny

OpenStudy (anonymous):

@jennybaeeee Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

wow no medals

OpenStudy (anonymous):

Plus your answer is wrong. @Jenn3251534 Multiply 3.14 * 9 * 9 @thedomainbunny

OpenStudy (anonymous):

Oops wrong tag @jennybaeeee *